OpenStudy (anonymous):

find the exact solution 100^x-1=1000^2x-1

OpenStudy (anonymous):

is that 100^(x-1) or 100^(x)-1?

OpenStudy (anonymous):

100^x-1

OpenStudy (anonymous):

your first 1 sorry

OpenStudy (anonymous):

100 = 10^2 1000 = 10^3 So 100^x-1 = 10^(2x-2) and 1000^2x-1 = 10^(6x-3) 2x-2 = 6x-3 4x = 1 x = 1/4

OpenStudy (anonymous):

the exact value of 1/4 is 0.25

OpenStudy (anonymous):

thank cortez

OpenStudy (anonymous):

you're welcome